OpenStudy (anonymous):

The perimeter of a square is 48 m. What is the length of a side and a diagonal?

OpenStudy (anonymous):

in a square all sides are equal so perimeter = 4 times length of side diagonal = length of side times sqrt2

OpenStudy (anonymous):

@myko it should be \[\sqrt{288}\]

OpenStudy (anonymous):

diagonal\[\sqrt{288}=12\sqrt{2}\]

OpenStudy (anonymous):

\[12\sqrt{2}\]